angles a and b are complementary angles, and angle b is 14degrees less than angle a. find the measures of angle a and angle b
Answer by lwsshak3(11628)

angles a and b are complementary angles, and angle b is 14degrees less than angle a. find the measures of angle a and angle b
**
a+b=90
b=a-14
a+a-14=90
2a=104
a=52
b=a-14=38
ans:
angle a=52º
angle b=38º